# Kalb-Ramond field, black holes and black strings in (2 + 1)D

### v4

@@ 3.1. New black string solutions
Note at $\gamma=0$ the metric describes BTZ black hole and it allows a non-zero Kalb-Ramond field strength $H$ [20]. The BTZ black hole solution (2.6) now with $\Phi=0$ and $B=B_{tx}dt\wedge dx$ where $B_{tx}=-\left[\rho^{2}-(\rho_{+}^{2}+\rho_{-}^{2})/2\right]/l$, i.e., the solution (3.15) at $\gamma=0$, is dual to the charged black string solution [20]14 In the appropriate gauge and with the appropriate coordinate changes., see also [[21]21,,  [22]22,,  [23]23,,  [24]24,,  [25]25] and [26]. We next consider different cases of the black string solution described by (3.15). We begin with the case where $\gamma$ is negative.15 The analytically extended solution, which is obtained by simply replacing $\rho$ by $i\widetilde{\rho}$ or $\rho^{2}$ by $r\in{\tenrm I\kern-1.79993ptR}$, in general contains closed time-like curves and thus we will not consider it. In the Penrose diagrams below we also do not include the regions which contain closed time-like curves.
@@ 3.2. Negative coupling
${\rho_{-}\over\rho_{+}}<\mu:={1+\gamma\over 1-\gamma}.$ (3.20)3.20( 3.20 )
Thus, in the case $\rho_{-}/\rho_{+}<\mu$ we have rotating black string solutions with inner and outer horizons but no curvature singularity. The solutions are exact solutions of classical string theory since they are related to $AdS_{3}$ primarily by T-dualities (and coordinate transformations). $\rho_{-}$ denotes the inner horizon and $\rho_{+}$ denotes the outer horizon. The period of the angle $x$ is chosen to avoid a conical singularity at the origin. We identify
$x\sim x+2\pi\sqrt{{\rho_{+}^{2}(1+\gamma)^{2}-\rho_{-}^{2}(1-\gamma)^{2}\over\rho^{2}_{+}-\rho^{2}_{-}}}.$ (3.21)3.21( 3.21 )
The global structure of the solutions which satisfy the inequality (3.20) is described using Penrose diagram in Fig. 1. The solutions are axially symmetric. Therefore, we set the (shifted) angular coordinate (in each patch) to a particular value, see Appendix D for details. Thus, a point in the diagram represents a circle in spacetime, i.e. the orbit of the axial isometry. The diagram extends infinitely in both vertical directions.
@@
$\rho_{0}={1\over 2}\sqrt{{\rho_{+}^{2}(1+\gamma)^{2}-\rho_{-}^{2}(1-\gamma)^{2}\over\gamma}}.$ (3.22)3.22( 3.22 )
@@
$\gamma={\rho_{-}-\rho_{+}\over\rho_{-}+\rho_{+}},\quad\rho_{+}>\rho_{-}.$ (3.23)3.23( 3.23 )
@@
$\eqalign{ds^{2}&={l^{2}\rho^{2}d\rho^{2}\over(\rho^{2}-\rho_{+}^{2})(\rho^{2}-\rho^{2}_{-})}-{\rho_{+}^{2}\over(1-\gamma)^{2}}{(\rho^{2}-\rho_{+}^{2})(\rho^{2}-\rho^{2}_{-})\over l^{2}\rho^{4}}dt^{2}+{\rho^{2}_{+}\over(1-\gamma)^{2}}\left(dx-{\rho_{+}\rho_{-}\over l\rho^{2}}dt\right)^{2},\cr B&=B_{tx}dt\wedge dx,\quad B_{tx}=-{1\over(1-\gamma)^{2}}{\rho_{+}^{2}\over l}\left(1-{\rho_{+}\rho_{-}\over\rho^{2}}\right),\quad Q=2\left({1-\gamma^{2}\over lg_{s}^{2}}\right),\cr e^{2\Phi}&=g_{s}^{2}|e^{2\phi}|={g_{s}^{2}\over(1-\gamma)^{2}}\left(\rho_{+}^{2}\over\rho^{2}\right).}$ (3.24)3.24( 3.24 )
The Ricci scalar is singular at the origin.16 Thus, here $x$ need not be periodic (with a definite period to avoid a conical singularity). Although we assumed $\rho_{+}>\rho_{-}$, i.e. $\gamma\neq 0$, to arrive at the solution, also here the limit $\rho_{+}\to\rho_{-}$ is mathematically well defined and gives non-trivial solution. In this limit, after an appropriate change of coordinates, we obtain $ds^{2}=(l^{2}/4)\left[dr^{2}/(r-r_{0})^{2}-d\tau^{2}(1-(r_{0}/r))^{2}+(d\chi-(r_{0}/r)d\tau)^{2}\right]$, $B=B_{\tau\chi}d\tau\wedge d\chi,B_{\tau\chi}=-(l^{2}/4)(1-(r_{0}/r))$, $e^{2\Phi}=g_{s}^{2}r_{0}/r$, where $r_{0}=\rho_{+}^{2}=\rho_{-}^{2}$. The scalar curvature is $R=\left[2(4r-7r_{0})r_{0}\right]/(l^{2}r^{2})$. Thus, $r=r_{0}$ is not a curvature singularity. We do not discuss this solution further here. We also note that both the dilaton and Kalb-Ramond fields are singular at the origin. We will discuss this family of solutions further later in the section. We in particular show that they are related to the solutions discussed in [21]. The global structure of the solutions for which $0\leq\rho_{0}<\rho_{-}$ is described using Penrose diagram in Fig. 2. The diagram is obtained in the same way as the example in Appendix D. In the diagram we set, in each patch, the (shifted) angular coordinate to a particular value since any other value is equivalent and related by axial symmetry. Therefore, a point in the diagram represents a circle in spacetime.
Fig. 2: The plot depicts the global structure of the black string solutions with $\rho_{-}\geq\mu\cdot\rho_{+}$, where $\mu=(1+\gamma)/(1-\gamma),-1<\gamma<0$. $\rho_{\pm}$ represent horizons. The solutions have a singularity at $\rho=\rho_{0}$. $\rho_{0}$ is represented by the zigzag lines. In general $0\leq\rho_{0}<\rho_{-}$. The singularity is time-like.
@@
$\eqalign{ds^{2}&={l^{2}\rho^{2}d\rho^{2}\over(\rho^{2}-\rho_{+}^{2})(\rho^{2}-\rho_{-}^{2})}-{1\over 4}{(\rho^{2}-\rho_{+}^{2})(\rho_{+}^{2}-\rho_{-}^{2})\over l^{2}\rho^{2}}dt^{2}+{1\over 4}{\rho^{2}(\rho_{+}^{2}-\rho_{-}^{2})\over\rho^{2}-\rho_{-}^{2}}\left(dx-{\rho_{+}\rho_{-}\over l\rho^{2}}dt\right)^{2},\cr B_{tx}&=-{1\over 4}\left({\rho_{+}^{2}-\rho_{-}^{2}\over l}\right)={\tenrm const.},\quad Q=0,\cr e^{2\Phi}&=g_{s}^{2}|e^{2\phi}|,\quad e^{-2\phi}=4\left({\rho^{2}-\rho_{-}^{2}\over\rho_{+}^{2}-\rho_{-}^{2}}\right).}$ (3.25)3.25( 3.25 )
@@
$R={4\over l^{2}}\left({\rho_{+}^{2}-\rho_{-}^{2}\over\rho^{2}-\rho_{-}^{2}}\right).$ (3.26)3.26( 3.26 )
Note the $B$ field, i.e. $B_{tx}$, is a non-zero constant. In general, using the gauge freedom we can set $B$ to zero. Thus, the solution describes a rotating dilaton black hole. This particular solution is already known in the literature [7]. It is obtained by directly solving the Einstein field equations with a scalar field. The novelty here, however, is that we simply obtained it using duality and coordinate transformations from global $AdS_{3}$ [1]. More on this black hole solution in the next section. The global structure of the solutions is described using Penrose diagram in Fig. 3. Each point represents a circle in spacetime.
@@
$\eqalign{ds^{2}&={l^{2}d\rho^{2}\over(\rho^{2}-\rho_{+}^{2})}-{\rho_{+}^{2}(\rho^{2}-\rho_{+}^{2})\over 4l^{2}\rho^{2}}dt^{2}+{1\over 4}\rho^{2}_{+}dx^{2},\cr B&=B_{tx}dt\wedge dx,\quad B_{tx}=-{1\over 4}{\rho_{+}^{2}\over l}={\tenrm const.},\quad Q=0,\cr e^{2\Phi}&={g_{s}^{2}\over 4}\left({\rho^{2}_{+}\over\rho^{2}}\right).}$ (3.27)3.27( 3.27 )
The curvature singularity is at the origin. This solution in the gauge in which $B=0$ is dual to the non-rotating BTZ black hole with a non-vanishing $B$ [20], see also [22].17 As we will show later in the paper, for a general gauge, it is also equivalent to a non-rotating black hole in a spacetime with non-vanishing angular velocity at radial infinity and no $B$ field. It is the product of the two dimensional black hole [[27]27,,  [28]28] with $S^{1}$. To see this define
${\rho\over\rho_{+}}={\tenrm cosh}\theta\geq 1.$ (3.28)3.28( 3.28 )
@@
$ds^{2}=l^{2}(d\theta^{2}-\tanh^{2}\theta d\tau^{2}+d\psi^{2}),$ (3.29)3.29( 3.29 )
@@
$\tau:={\rho_{+}\over 2l}t,\quad\psi:={\rho_{+}\over 2}x.$ (3.30)3.30( 3.30 )
@@
$e^{-2\Phi}={4\over g_{s}^{2}}{\tenrm cosh}^{2}\theta.$ (3.31)3.31( 3.31 )
@@
$R={4\over l^{2}}{\tenrm sech}^{2}\theta.$ (3.32)3.32( 3.32 )

### v5

@@ 3.1. New black string solutions
We can always put the solution (3.15), for $\gamma\neq 0$, into the form
$\eqalign{ds^{2}&=L^{-2}(\rho)d\rho^{2}-N^{2}(\rho)d{\widehat{t}}^{2}+{(\rho^{2}_{+}-\rho^{2}_{-})\over 4|\gamma|}\left(d{\widehat{x}}+Zd{\widehat{t}}\right)^{2},\cr B&=B_{{\widehat{t}}{\widehat{x}}}d{\widehat{t}}\wedge d{\widehat{x}},\quad B_{{\widehat{t}}{\widehat{x}}}=\left({\rho_{+}^{2}-\rho_{-}^{2}\over 2l}\right)\cdot\left[{\rho_{+}^{2}(1+\gamma)+\rho_{-}^{2}(1-\gamma)-2\rho^{2}\over\rho_{+}^{2}(1+\gamma)^{2}-\rho_{-}^{2}(1-\gamma)^{2}-4\gamma\rho^{2}}\right],\cr e^{2\Phi}&=g_{s}^{2}\left|{\rho_{+}^{2}-\rho_{-}^{2}\over\rho_{+}^{2}(1+\gamma)^{2}-\rho_{-}^{2}(1-\gamma)^{2}-4\gamma\rho^{2}}\right|,}$ (3.20)3.20( 3.20 )
by unwrapping $x$ and applying a coordinate transformation. The metric components, i.e., $L$, $N$ and $Z$, are
$\eqalign{L^{2}&={(\rho^{2}-\rho_{-}^{2})(\rho^{2}-\rho_{+}^{2})\over l^{2}\rho^{2}},\cr N^{2}&={4|\gamma|(\rho_{+}^{2}-\rho_{-}^{2})\over l^{2}}\cdot{(\rho^{2}-\rho_{-}^{2})(\rho^{2}-\rho_{+}^{2})\over[\rho_{+}^{2}(1+\gamma)^{2}-\rho_{-}^{2}(1-\gamma)^{2}-4\gamma\rho^{2}]^{2}},\cr Z&=\left(\gamma\over|\gamma|\right)\cdot\left(1-\gamma^{2}\over l\right)\cdot\left[{\rho_{+}^{2}-\rho_{-}^{2}\over\rho_{+}^{2}(1+\gamma)^{2}-\rho_{-}^{2}(1-\gamma)^{2}-4\gamma\rho^{2}}\right].}$ (3.21)3.21( 3.21 )
For $\gamma<0$, the coordinates are related by the Lorentz boost $t={\widehat{t}}{\tenrm cosh}(\xi)+l{\widehat{x}}{\tenrm sinh}(\xi),lx=l{\widehat{x}}{\tenrm cosh}(\xi)+{\widehat{t}}{\tenrm sinh}(\xi)$, where $\xi=(1/2)\cdot\ln\{|\gamma|\cdot\left[(\rho_{+}+\rho_{-})/(\rho_{+}-\rho_{-})\right]\}$. Alternatively, we can take $\xi=(1/2)\cdot\ln\{(1/|\gamma|)\cdot\left[(\rho_{+}+\rho_{-})/(\rho_{+}-\rho_{-})\right]\}$. The latter choice only changes $Z\to-Z$. Note the $B$ field is invariant under the boost. Thus, also the axion charge $Q$ is invariant. For $\gamma>0$, the coordinates are instead related by the (general) coordinate transformation $lx=-{\widehat{t}}{\tenrm cosh}(\xi)+l{\widehat{x}}{\tenrm sinh}(\xi),t=l{\widehat{x}}{\tenrm cosh}(\xi)-{\widehat{t}}{\tenrm sinh}(\xi)$, where $\xi$ is as given above.14 Note the (general) coordinate transformation, for positive $\gamma$, is not a Lorentz (boost) transformation. It is a combination of a Lorentz boost and a spatial-temporal rotation (of $\pi/2$). We obtain the boost transformation by making a spacetime rotation, i.e., the replacements ${\widehat{x}}\to{\widehat{t}}/l,{\widehat{t}}\to-l{\widehat{x}}$. Also for $\gamma>0$, the latter choice for $\xi$ only changes $Z\to-Z$. The $B$ field, also for $\gamma>0$, is invariant under the (general) coordinate transformation.
The black hole solution (3.20) extends (or winds around) and, for non-constant $Z$, also rotates along the (compact) ${\widehat{x}}$ direction. For this reason, therefore, we refer to the solution (3.15) as rotating black string. When $Z$ is a constant, we simply redefine ${\widehat{x}}$ as ${\widehat{x}}\to{\widehat{x}}-Z{\widehat{t}}$ to effectively set $Z$ to zero. Therefore, for $Z=0$ (or, more generally, for constant $Z$), the black hole is a trivial bundle over (the spatial dimension) ${\widehat{x}}$ and, it is commonly referred to as simply black string. Note that a black string does not always necessarily contain a black hole factor. That is, the fiber may not describe or solve a black hole solution. However, in a boosted and rotated frame or in a different coordinate system it could describe a black hole. A simple example is the solution at $\gamma=-1$. At $\gamma=-1$, we note that $Z=0$. The solution has the product form ${\cal M}_{2}\times{\tenrm I\kern-1.79993ptR}$. The factor ${\cal M}_{2}$ does not describe a black hole in two dimensions. However, as we momentarily show, we can define a new coordinate system in which the solution contains a two dimensional black hole factor. Also, a product spacetime of a black hole solution and a line ${\tenrm I\kern-1.79993ptR}$ or a circle $S^{1}$ does not always necessarily solve the string theory equations of motion and thus, it does not always necessarily describe or lead to a black string solution [22].
The solution (3.20) is related to the black string solution [23]. However, the relation is not direct. We explain this now. To this end, it is convenient to introduce (assuming $\gamma\neq 0$) the new coordinate ${\widehat{\rho}}$,
$4|\gamma|{\widehat{\rho}}^{2}=-4\gamma\rho^{2}+\left[\rho_{+}^{2}(1+\gamma)^{2}-\rho_{-}^{2}(1-\gamma)^{2}\right].$ (3.22)3.22( 3.22 )
Note that by shifting (and/or analytically continuing) $\rho^{2}$ we are effectively fixing $|\gamma|$ to be equal to $(\rho_{+}-\rho_{-})/(\rho_{+}+\rho_{-})$. This will become more clear later in the section. In terms of the new coordinates $({\widehat{t}},{\widehat{x}},{\widehat{\rho}})$ the solution (3.20) takes the form
$\eqalign{ds^{2}&={l^{2}{\widehat{\rho}}^{2}d{\widehat{\rho}}^{2}\over({\widehat{\rho}}^{2}-{\widehat{\rho}}_{-}^{2})({\widehat{\rho}}^{2}-{\widehat{\rho}}_{+}^{2})}-{{\widehat{\rho}}^{2}_{+}\over\left(1+|\gamma|\right)^{2}}\left[({\widehat{\rho}}^{2}-{\widehat{\rho}}^{2}_{+})({\widehat{\rho}}^{2}-{\widehat{\rho}}^{2}_{-})\over l^{2}{\widehat{\rho}}^{4}\right]d{\widehat{t}}^{2}+{{\widehat{\rho}}^{2}_{+}\over\left(1+|\gamma|\right)^{2}}\left[d{\widehat{x}}+\left({\gamma\over|\gamma|}\right){{\widehat{\rho}}_{-}{\widehat{\rho}}_{+}\over l{\widehat{\rho}}^{2}}d{\widehat{t}}\right]^{2},\cr B&=B_{{\widehat{t}}{\widehat{x}}}d{\widehat{t}}\wedge d{\widehat{x}},\quad B_{{\widehat{t}}{\widehat{x}}}=\left({|\gamma|\over\gamma}\right)\cdot{1\over\left(1+|\gamma|\right)^{2}}\cdot{{\widehat{\rho}}^{2}_{+}\over l}\cdot\left[1-{{\widehat{\rho}}_{+}{\widehat{\rho}}_{-}\over{\widehat{\rho}}^{2}}\right],\cr e^{2\Phi}&={g_{s}^{2}\over\left(1+|\gamma|\right)^{2}}\left({{\widehat{\rho}}_{+}^{2}\over{\widehat{\rho}}^{2}}\right),}$ (3.23)3.23( 3.23 )
where ${\widehat{\rho}}_{+}$ and ${\widehat{\rho}}_{-}$ are given in terms of $\rho_{+}$, $\rho_{-}$ and $\gamma$,
${\widehat{\rho}}^{2}_{+}={(1+|\gamma|)^{2}\over 4|\gamma|}(\rho_{+}^{2}-\rho_{-}^{2}),\quad{\widehat{\rho}}^{2}_{-}={(1-|\gamma|)^{2}\over 4|\gamma|}(\rho_{+}^{2}-\rho_{-}^{2}).$ (3.24)3.24( 3.24 )
Note we have effectively avoided the need for the absolute value sign in (3.4) when $\gamma>0$ by shifting $\rho$ (3.22), or equivalently, by restricting $\rho$ to the region $\rho<\rho_{+}$. We have also shrunk/moved the curvature singularity to the origin. Moreover, we can eliminate the explicit dependence of (3.23) on $\gamma$ by redefining $g_{s}$ and the two coordinates ${\widehat{t}}$ and ${\widehat{x}}$. Therefore, the dependence on $\gamma$ is implicit. To obtain this implicit dependence we take a closer look at (3.24). The relations (3.24) imply the constraint15 Upon an appropriate change of coordinates, the solution (3.23) becomes $\eqalign{ds^{2}&={l^{2}\over 4}\left[{dr^{2}\over(r-r_{-})(r-r_{+})}-{(r-r_{-})(r-r_{+})\over r^{2}}d\tau^{2}+\left(d\chi+{\gamma\over|\gamma|}{\sqrt{r_{-}r_{+}}\over r}d\tau\right)^{2}\right],\cr B&=B_{\tau\chi}d\tau\wedge d\chi,\quad B_{\tau\chi}={l^{2}\over 4}\left[1-{\sqrt{r_{-}r_{+}}\over r}\right],\cr\Phi&={1\over 2}\ln\left(r_{+}\over r\right)+\iota,\quad\iota=\ln\left({g_{s}\over 1+|\gamma|}\right),\quad{r_{-}\over r_{+}}=\left({1-|\gamma|\over 1+|\gamma|}\right)^{2}.}$ (3.25)3.25( 3.25 ) The scalar curvature is $R={2\over l^{2}r^{2}}\left[2(r_{+}+r_{-})r-7r_{+}r_{-}\right]$. Thus, it is singular at the origin $r=0$.
${{\widehat{\rho}}_{-}\over{\widehat{\rho}}_{+}}={1-|\gamma|\over 1+|\gamma|},\quad|\gamma|={{\widehat{\rho}}_{+}-{\widehat{\rho}}_{-}\over{\widehat{\rho}}_{+}+{\widehat{\rho}}_{-}}.$ (3.26)3.26( 3.26 )
Thus, the ratio ${\widehat{\rho}}_{-}/{\widehat{\rho}}_{+}$ is fixed by $\gamma$. In particular, at $\gamma=-1$ we have ${\widehat{\rho}}_{-}=0$. Therefore, in the new coordinates, the solution at $\gamma=-1$ has the product form ${\cal M}_{2}\times{\tenrm I\kern-1.79993ptR}$ where ${\cal M}_{2}$ is a black hole [[24]24,,  [25]25]. Later, we will explicitly show that (3.23) is related indirectly to the non-rotating black string solution [23] by a further Lorentz boost.16 In (3.23) we could treat ${\widehat{\rho}}_{\pm}$ more generally as independent parameters and simply ignore their dependence on $\gamma$ (3.24) or the constraint (3.26). However, this is a separate case. To the best of my knowledge the black string solution (3.15) has not been reported elsewhere in the literature.
@@
Note at $\gamma=0$ the metric describes BTZ black hole and it allows a non-zero Kalb-Ramond field strength $H$ [26].17 Therefore, one may as well prefer to refer to the solutions (labelled by the triplet $(\rho_{-},\rho_{+};\gamma)$) (3.15) as deformed black holes. The BTZ black hole solution (2.6) now with $\Phi=0$ and $B=B_{tx}dt\wedge dx$ where $B_{tx}=-\left[\rho^{2}-(\rho_{+}^{2}+\rho_{-}^{2})/2\right]/l$, i.e., the solution (3.15) at $\gamma=0$, is dual to the charged black string solution [26]18 In the appropriate gauge and with the appropriate coordinate changes., see also [[23]23,,  [27]27,,  [28]28,,  [29]29,,  [30]30] and [31]. We next consider different cases of the black string solution described by (3.15). We begin with the case where $\gamma$ is negative.19 The analytically extended solution, which is obtained by simply replacing $\rho$ by $i\widetilde{\rho}$ or $\rho^{2}$ by $r\in{\tenrm I\kern-1.79993ptR}$, in general contains closed time-like curves and thus we will not consider it. In the Penrose diagrams below we also do not include the regions which contain closed time-like curves.
@@ 3.2. Negative coupling
${\rho_{-}\over\rho_{+}}<\mu:={1+\gamma\over 1-\gamma}.$ (3.27)3.27( 3.27 )
Thus, in the case $\rho_{-}/\rho_{+}<\mu$ we have rotating black string solutions with inner and outer horizons but no curvature singularity. The solutions are exact solutions of classical string theory since they are related to ${\cal A}_{3}$ primarily by coordinate transformations. $\rho_{-}$ denotes the inner horizon and $\rho_{+}$ denotes the outer horizon. The period of the angle $x$ is chosen to avoid a conical singularity at the origin. We identify
$x\sim x+2\pi\sqrt{{\rho_{+}^{2}(1+\gamma)^{2}-\rho_{-}^{2}(1-\gamma)^{2}\over\rho^{2}_{+}-\rho^{2}_{-}}}.$ (3.28)3.28( 3.28 )
The global structure of the solutions which satisfy the inequality (3.27) is described using Penrose diagram in Fig. 1. The solutions are axially symmetric. Therefore, we set the (shifted) angular coordinate (in each patch) to a particular value, see Appendix D for details. Thus, a point in the diagram represents a circle in spacetime, i.e. the orbit of the axial symmetry. The diagram extends infinitely in both vertical directions.
@@
$\rho_{0}={1\over 2}\sqrt{{\rho_{+}^{2}(1+\gamma)^{2}-\rho_{-}^{2}(1-\gamma)^{2}\over\gamma}}.$ (3.29)3.29( 3.29 )
@@
$\gamma={\rho_{-}-\rho_{+}\over\rho_{-}+\rho_{+}},\quad\rho_{+}>\rho_{-}.$ (3.30)3.30( 3.30 )
@@
$\eqalign{ds^{2}&={l^{2}\rho^{2}d\rho^{2}\over(\rho^{2}-\rho_{+}^{2})(\rho^{2}-\rho^{2}_{-})}-{\rho_{+}^{2}\over(1-\gamma)^{2}}{(\rho^{2}-\rho_{+}^{2})(\rho^{2}-\rho^{2}_{-})\over l^{2}\rho^{4}}dt^{2}+{\rho^{2}_{+}\over(1-\gamma)^{2}}\left(dx-{\rho_{+}\rho_{-}\over l\rho^{2}}dt\right)^{2},\cr B&=B_{tx}dt\wedge dx,\quad B_{tx}=-{1\over(1-\gamma)^{2}}{\rho_{+}^{2}\over l}\left(1-{\rho_{+}\rho_{-}\over\rho^{2}}\right),\quad Q=2\left({1-\gamma^{2}\over lg_{s}^{2}}\right),\cr e^{2\Phi}&=g_{s}^{2}|e^{2\phi}|={g_{s}^{2}\over(1-\gamma)^{2}}\left(\rho_{+}^{2}\over\rho^{2}\right).}$ (3.31)3.31( 3.31 )
We first note that (3.31) is the same as (3.23). The Ricci scalar is singular at the origin.20 Thus, here $x$ need not be periodic (with a definite period to avoid a conical singularity). Although we assumed $\rho_{+}>\rho_{-}$, i.e. $\gamma\neq 0$, to arrive at the solution, also here the limit $\rho_{+}\to\rho_{-}$ is mathematically well-defined and gives a non-trivial solution. In this limit, after an appropriate change of coordinates, we obtain $ds^{2}=(l^{2}/4)\left[dr^{2}/\left(r-r_{0}\right)^{2}-d\tau^{2}\left(1-(r_{0}/r)\right)^{2}+\left(d\chi-(r_{0}/r)d\tau\right)^{2}\right]$, $B=B_{\tau\chi}d\tau\wedge d\chi$, $B_{\tau\chi}=-(l^{2}/4)\left[1-(r_{0}/r)\right]$, $e^{2\Phi}=g_{s}^{2}r_{0}/r$, where $r_{0}=\rho_{+}^{2}=\rho_{-}^{2}$. The scalar curvature is $R=\left[2(4r-7r_{0})r_{0}\right]/(l^{2}r^{2})$. Thus, $r=r_{0}$ is not a curvature singularity. We do not discuss this solution further here. We also note that both the dilaton and Kalb-Ramond fields are singular at the origin. We will discuss this family of solutions further later in the section. We in particular show that they are related indirectly to the solutions discussed in [23]. The global structure of the solutions for which $0\leq\rho_{0}<\rho_{-}$ is described using Penrose diagram in Fig. 2. The diagram is obtained in the same way as the example in Appendix D. In the diagram we set, in each patch, the (shifted) angular coordinate to a particular value since any other value is equivalent and related by axial symmetry. Therefore, a point in the diagram represents a circle in spacetime.
Fig. 2: The plot depicts the global structure of the black string solutions with $\rho_{-}\geq\mu\cdot\rho_{+}$, where $\mu=(1+\gamma)/(1-\gamma),-1<\gamma<0$. $\rho_{\pm}$ represent horizons. The solutions have a singularity at $\rho=\rho_{0}$. $\rho_{0}$ is represented by the zigzag lines. In general $0\leq\rho_{0}<\rho_{-}$. The singularity is time-like. For $\rho_{0}=0$, the maximally extended solution has no closed time-like curves. Therefore, for $\rho_{0}=0$, the diagram can be extended behind $\rho=\rho_{0}=0$.
@@
$\eqalign{ds^{2}&={l^{2}\rho^{2}d\rho^{2}\over(\rho^{2}-\rho_{+}^{2})(\rho^{2}-\rho_{-}^{2})}-{1\over 4}{(\rho^{2}-\rho_{+}^{2})(\rho_{+}^{2}-\rho_{-}^{2})\over l^{2}\rho^{2}}dt^{2}+{1\over 4}{\rho^{2}(\rho_{+}^{2}-\rho_{-}^{2})\over\rho^{2}-\rho_{-}^{2}}\left(dx-{\rho_{+}\rho_{-}\over l\rho^{2}}dt\right)^{2},\cr B_{tx}&=-{1\over 4}\left({\rho_{+}^{2}-\rho_{-}^{2}\over l}\right)={\tenrm const.},\quad Q=0,\cr e^{2\Phi}&=g_{s}^{2}|e^{2\phi}|,\quad e^{-2\phi}=4\left({\rho^{2}-\rho_{-}^{2}\over\rho_{+}^{2}-\rho_{-}^{2}}\right).}$ (3.32)3.32( 3.32 )
@@
$R={4\over l^{2}}\left({\rho_{+}^{2}-\rho_{-}^{2}\over\rho^{2}-\rho_{-}^{2}}\right).$ (3.33)3.33( 3.33 )
Note the $B$ field, i.e. $B_{tx}$, is a non-zero constant. In general, using the gauge freedom we can set $B$ to zero. Thus, the solution describes a rotating dilaton black hole. This particular solution is already known in the literature [7]. It is obtained by directly solving the Einstein field equations with a minimally coupled scalar field. The novelty here, however, is that we simply obtained it using duality and coordinate transformations from global $AdS_{3}$ [1]. More on this black hole solution in the next section. The global structure of the solutions is described using Penrose diagram in Fig. 3. Each point represents a circle in spacetime.
@@
$\eqalign{ds^{2}&={l^{2}d\rho^{2}\over(\rho^{2}-\rho_{+}^{2})}-{\rho_{+}^{2}(\rho^{2}-\rho_{+}^{2})\over 4l^{2}\rho^{2}}dt^{2}+{1\over 4}\rho^{2}_{+}dx^{2},\cr B&=B_{tx}dt\wedge dx,\quad B_{tx}=-{1\over 4}{\rho_{+}^{2}\over l}={\tenrm const.},\quad Q=0,\cr e^{2\Phi}&={g_{s}^{2}\over 4}\left({\rho^{2}_{+}\over\rho^{2}}\right).}$ (3.34)3.34( 3.34 )
The curvature singularity is at the origin. This solution in the gauge in which $B=0$ is dual to the non-rotating BTZ black hole with a non-vanishing $B$ [26].21 As we will show later in the paper, for a general gauge, it is also equivalent to a non-rotating black hole in a spacetime with non-vanishing angular velocity at radial infinity and no $B$ field. It is also related indirectly to [23], see [27]. It is the product of the two dimensional black hole [[24]24,,  [25]25] with $S^{1}$. To see this define
${\rho\over\rho_{+}}={\tenrm cosh}\theta\geq 1.$ (3.35)3.35( 3.35 )
@@
$ds^{2}=l^{2}(d\theta^{2}-\tanh^{2}\theta d\tau^{2}+d\psi^{2}),$ (3.36)3.36( 3.36 )
@@
$\tau:={\rho_{+}\over 2l}t,\quad\psi:={\rho_{+}\over 2}x.$ (3.37)3.37( 3.37 )
@@
$e^{-2\Phi}={4\over g_{s}^{2}}{\tenrm cosh}^{2}\theta.$ (3.38)3.38( 3.38 )
@@
$R={4\over l^{2}}{\tenrm sech}^{2}\theta.$ (3.39)3.39( 3.39 )